# Context-Alignment: Activating and Enhancing LLM Capabilities in Time Series

### v1

@@ 5 Conclusion
In this paper, We point out that effectively utilizing LLMs in TS tasks requires first activating their capabilities, then enhancing their performance. By rethinking the inherent strength of LLMs, We are the first to propose Context-Alignment paradigm, which aims to construct a context-level alignment between TS and language. Unlike previous methods based on token-level alignment, Context-Alignment constructs a consistent context for TS-language multimodal inputs, better harnessing LLMs’ deep understanding on context to activate their potential on TS tasks. We develop a framework, DSCA-GNNs, to achieve Context-Alignment. Besides, by integrating the Demonstration Examples Prompt technique, we introduce DECA, which enhances LLMs’ performance in TS tasks. Experiments demonstrate our method significantly outperforms others, particularly in few-shot and zero-shot forecasting tasks, and ablation studies confirm the importance of Context-Alignment.
@@ A.1 IMPLEMENTATION
All deep learning networks are implemented in PyTorch and trained on NVIDIA H800 80GB GPUs and GeForce RTX 4090 GPUs. We conduct our experiments using the pretrained models from Wolf et al. [2020]. To ensure fair comparisons, we adhere to the experimental configurations outlined in Wu et al. [2022] across all baselines and maintain a uniform evaluation procedure. As discussed in the ablation study in Sec. 4.7 (The Number of LLM Layers), we adopt the first 4 layers of GPT-2. For other LLM-based methods Zhou et al. [2023], Jin et al. [2024], Pan et al. [2024], we strictly follow their provided experimental settings or cite their performance if applicable. Most of our proposed method’s training settings are based on Zhou et al. [2023]: For predictive tasks, we utilize DECA with $N=2$ in forms forms 5 and 6, providing one demonstration example. The Adam optimizer is used with decay rates $\beta=(0.9,0.999)$ and initial learning rates from $\{10^{-4},5\times 10^{-4}\}$. We implement a cosine annealing schedule with $T_{\text{max}}=20$ and $\eta_{\text{min}}=10^{-8}$, and set the batch size to 256. Early stopping is configured throughout the training process. MSE loss is employed for long-term forecasting, while SMAPE loss is used for short-term predictions. In classification tasks, as described in Sec. 4.6, the DECA framework is applied to binary class datasets, including a demonstration example for each category. For multi-class datasets, such as handwriting recognition with 26 classes, we employ the VCA approach due to the input constraints of LLMs. The RAdam optimizer with initial learning rates from $\{10^{-2},10^{-3}\}$ and a batch size of 64 is used. Training also incorporates early stopping and employs cross-entropy loss.
@@ A.2 DATASET DETAILS
For the long-term forecasting task, we utilized eight widely used multivariate datasets Wu et al. [2022], as detailed in Table 7. These include the Electricity Transformer Temperature (ETT) datasets Zhou et al. [2021], as well as Illness, Weather, Electricity, and Traffic datasets. The ETT dataset comprises ETTh1 and ETTh2, while the ETTm dataset includes ETTm1 and ETTm2. Specifically, the ETT datasets contain power load data from two power stations at varying resolutions; the Weather dataset features 21 meteorological indicators from Germany; the ILI dataset captures weekly patient counts and influenza-like illness rates; the Electricity dataset consists of hourly consumption data from 321 customers; and the Traffic dataset records road occupancy from various sensors on San Francisco highways. Table 7 consolidates the feature details of these datasets, with the ETT dataset being applied in both few-shot and zero-shot learning tasks.
@@
For the short-term forecasting task, we employed the M4 benchmark dataset Makridakis et al. [2018], which includes 10,000 time series across various domains, from business to economic forecasting, as shown in Table 8. The time series data is categorized into six groups, with sampling rates ranging from annually to hourly.
@@ A.3 BASELINE DETAILS
iTransformer Liu et al. [2023a]: iTransformer repurposes the Transformer architecture by applying attention and feed-forward networks on inverted dimensions to improve multivariate time series forecasting, achieving state-of-the-art performance on real-world datasets.
@@
FEDformer Zhou et al. [2022]: FEDformer combines Transformer models with seasonal-trend decomposition to capture both global trends and detailed structures in time series data.
@@
Stationary Liu et al. [2022]: Stationary introduces Non-stationary Transformers, which include Series Stationarization for improved predictability and De-stationary Attention to restore intrinsic non-stationary information, resulting in significant performance improvements and a substantial reduction in mean squared error compared to mainstream Transformer models.
@@
ETSFormer Woo et al. [2022]: ETSFormer is a novel Transformer architecture designed specifically for time-series forecasting, addressing the limitations of traditional models by incorporating exponential smoothing principles.
@@
PatchTST Nie et al. [2022]: PatchTST introduces an efficient Transformer-based architecture for multivariate time series forecasting and self-supervised representation learning, utilizing a segmentation approach that divides time series into subseries-level patches. This design enhances local semantic retention, significantly reduces computation and memory usage of attention maps, and allows the model to consider longer historical data.
@@
TimesNet Wu et al. [2022]: TimesNet is introduced as a novel approach for time series analysis that focuses on modeling temporal variations by transforming 1D time series into 2D tensors, thereby capturing complex intraperiod and interperiod variations.
@@
Dlinear Zeng et al. [2023]: Dlinear challenges the efficacy of Transformer-based models for long-term time series forecasting (LTSF), highlighting their limitations in capturing temporal relationships due to the permutation-invariant nature of self-attention mechanisms.
@@
N-HiTs Challu et al. [2023]: N-HiTs introduces a novel model for long-horizon forecasting that utilizes hierarchical interpolation and multi-rate data sampling techniques to effectively address the challenges of prediction volatility and computational complexity.
@@
N-BEATS Oreshkin et al. [2019]: N-BEATS addresses the univariate time series point forecasting problem using a novel deep neural architecture featuring backward and forward residual links with a deep stack of fully connected layers. Importantly, the model’s configuration without time-series-specific components suggests that deep learning primitives alone can effectively tackle a variety of forecasting challenges while also providing interpretable outputs with minimal accuracy loss.
@@
GPT4TS Zhou et al. [2023]: GPT4TS presents a novel approach that leverages pre-trained LLMs for general time series analysis, addressing the challenge of limited training data by utilizing the Frozen Pretrained Transformer (FPT) architecture without altering the self-attention and feedforward layers.
@@
Time-LLM Jin et al. [2024]: Time-LLM is a reprogramming framework designed to adapt LLMs for general time series forecasting by aligning time series data with natural language through input transformation and context enrichment techniques. Here, we use GPT-2 Radford et al. [2019] as the base LLM.
@@
S2IP-LLM Pan et al. [2024]: S2IP-LLM aligns pre-trained LLMs with time series embeddings to enhance forecasting performance. By designing a tokenization module for cross-modality alignment and utilizing semantic anchors from pre-trained word embeddings, S2IP-LLM effectively encodes temporal dynamics and retrieves relevant context for time series.
@@ A.4 Evaluation Metrics
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of long-term, few-shot, and zero-shot forecasting. For short-term forecasting on the M4 benchmark, following the approach in N-BEATS Oreshkin et al. [2019], we adopt Symmetric Mean Absolute Percentage Error (SMAPE), Mean Absolute Scaled Error (MASE), and Overall Weighted Average (OWA). For time series classification tasks, as referenced in GPT4TS Zhou et al. [2023], accuracy is used as the evaluation metric.
@@
where $H$ denote the number of data samples, corresponding to the forecasting horizon in the experiment. $s$ represents the periodicity of the time series. $Y_{h}$ and $\hat{Y}_{h}$ refer to the $h$-th ground truth and its corresponding prediction, respectively.
@@ Appendix B Dual-Scale Context-Alignment GNNs in Classification
The construction of Dual-Scale Context-Alignment GNNs relies on the given prompt. In this section, we introduce the Dual-Scale Context-Alignment GNNs based on vanilla prompt and demonstration examples prompt in classification tasks. Due to input length constraints, it is challenging to provide examples for many categories. Therefore, for multi-category classification tasks, we employ VCA approach. For binary classification tasks, we use DECA and select one example from each category to serve as fixed input examples for both training and testing.
@@ B.1 Vanilla Context-Alignment (VCA)
We only replace the prompt content in Sec. 3.2 with “Predict category ($x$ in total) using previous data:”, and the method for constructing the graph remains similar as Sec. 3.2, where $x$ denotes the number of categories.
B.2 Demonstration Examples Based Context-Alignment (DECA)
@@
$\{{\bm{z}}_{1}^{(k)},\dots,{\bm{z}}_{m}^{(k)}\}=\{{\bm{z}}_{1}^{(l+1)},\dots,{\bm{z}}_{m}^{(l+1)}\}=\{{\bm{z}}_{1},\dots,{\bm{z}}_{m}\}$ are the token embeddings of the prompt. We utilise $\{{\bm{e}}_{1}^{(k)},\dots,{\bm{e}}_{n}^{(k)}\}$ as fixed examples during training and testing phases, where $\{{\bm{e}}_{1}^{(k)},\dots,{\bm{e}}_{n}^{(k)}\}$ is from the $k$-th categories and $k=1,\dots,l$. ${\bm{y}}^{(k)}$ is the embedding of the correct label for $\{{\bm{e}}_{1}^{(k)},\dots,{\bm{e}}_{n}^{(k)}\}$. $\{{\bm{e}}_{1}^{(l+1)},\dots,{\bm{e}}_{n}^{(l+1)}\}$ is the TS that needs to be classified. In the fine-grained GNN $G_{F}$, each element in the form 10 is treated as a node. Similar with Sec. 3.2 and Sec. 3.3, two learnable linear layers $f_{e}$, $f_{z}$ map form 10 to form 11. Every element in form 11 belongs to an $M$-dimensional space.
@@
The directed edges of $G_{F}$ are decompositions of the directed edges of $G_{C}$, which can be represented by formula LABEL:edgeFC.
@@
Similar to the demo in Sec. 3.2, we introduce an assignment matrix and a learnable weight matrix to achieve learnable interactions between the two scales. Both $\hat{\bm{N}}_{F}$ and $\hat{\bm{N}}_{C}$ are input into pre-trained LLMs. The Dual-Scale Context-Alignment GNNs can be flexibly integrated into various layers of pre-trained LLMs as depicted in Fig. 1, and only the first time apply it need $f_{e}$ and $f_{z}$ to obtain coarse-grained GNN. The output from the $G_{F}$ branch is used to compute the cross entropy loss against the ground truth.
@@ C.1 Long-term Forecasting full results
Table 10 presents the detailed long-term forecasting results across four prediction horizons. Compared to other models, including LLM-based methods, Transformer-based models, and other high-performing approaches, DECA demonstrates strong and relatively stable performance across various datasets. It achieves an average 3.1% MSE reduction compared to the second-best method, PatchTST, and outperforms LLM-based methods (S2IP-LLM, Time-LLM, and GPT4TS) by 7.3%, 12.2%, and 16.6%, respectively. We attribute this consistent advantage to the successful Context-Alignment, which effectively guides LLMs’ deep understanding of time series data, with demonstration examples prompt serving as a key factor in enhancement.
@@
width = 1colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,18,23,28,33,381 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18,23,28,33,38 = 1-240.05em, hline43 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21,24-26,29-31,34-36,39-41, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18,23,28,33,38abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22,27,32,37,42belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer

### v2

@@ 5 Conclusion
In this paper, we point out that effectively utilizing LLMs in TS tasks requires first activating their capabilities, and then enhancing their performance. By rethinking the inherent strength of LLMs, we are the first to propose Context-Alignment paradigm, which aims to construct a context-level alignment between TS and language. Unlike previous methods based on token-level alignment, Context-Alignment constructs a consistent context for TS-language multimodal inputs, better harnessing LLMs’ deep understanding of context to activate their potential on TS tasks. We develop DSCA-GNNs to achieve Context-Alignment. Besides, by integrating the Demonstration Examples Prompt technique, we introduce FSCA, which enhances LLMs’ performance in TS tasks. Experiments demonstrate that our method significantly outperforms others, particularly in few-shot and zero-shot forecasting tasks, and ablation studies confirm the importance of Context-Alignment.
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (Grant No. 62106116), the China Meteorological Administration (Grant No. QBZ202316), the National Key Research and Development Program of China (Grant No. 2024YFF1500600), and the Major Science and Technology Projects of Ningbo (Grant No. 2022Z236), as well as by the High Performance Computing Centers at Eastern Institute of Technology, Ningbo, and Ningbo Institute of Digital Twin.
@@ A.1 IMPLEMENTATION
All deep learning networks are implemented in PyTorch and trained on NVIDIA H800 80GB GPUs and GeForce RTX 4090 GPUs. We conduct our experiments using the pre-trained models from Wolf et al. [2020]. To ensure fair comparisons, we adhere to the experimental configurations outlined in Wu et al. [2022] across all baselines and maintain a uniform evaluation procedure. As discussed in the ablation study in Sec. 4.7 (The Number of LLM Layers), we adopt the first 4 layers of GPT-2. For other LLM-based methods [Zhou et al., 2023, Jin et al., 2024, Pan et al., 2024], we strictly follow their provided experimental settings or cite their performance if applicable. Most of our proposed method’s training settings are based on Zhou et al. [2023]: For predictive tasks, we utilize FSCA with $N=2$ in forms 5 and 6, providing one demonstration example. The Adam optimizer is used with decay rates $\beta=(0.9,0.999)$ and initial learning rates from $\{10^{-4},5\times 10^{-4}\}$. We implement a cosine annealing schedule with $T_{\text{max}}=20$ and $\eta_{\text{min}}=10^{-8}$, and set the batch size to 256. Early stopping is configured throughout the training process. MSE loss is employed for long-term forecasting, while SMAPE loss is used for short-term predictions. In classification tasks, as described in Sec. 4.6, the FSCA framework is applied to binary class datasets, including a demonstration example for each category. For multi-class datasets, such as handwriting recognition with 26 classes, we employ the VCA approach due to the input constraints of LLMs. The RAdam optimizer with initial learning rates from $\{10^{-2},10^{-3}\}$ and a batch size of 64 is used. Training also incorporates early stopping and employs cross-entropy loss.
@@ A.2 DATASET DETAILS
For the long-term forecasting task, we utilized eight widely used multivariate datasets [Wu et al., 2022], as detailed in Table 7. These include the Electricity Transformer Temperature (ETT) datasets [Zhou et al., 2021], as well as Illness, Weather, Electricity, and Traffic datasets. The ETT dataset comprises ETTh1 and ETTh2, while the ETTm dataset includes ETTm1 and ETTm2. Specifically, the ETT datasets contain power load data from two power stations at varying resolutions; the Weather dataset features 21 meteorological indicators from Germany; the ILI dataset captures weekly patient counts and influenza-like illness rates; the Electricity dataset consists of hourly consumption data from 321 customers; and the Traffic dataset records road occupancy from various sensors on San Francisco highways. Table 7 consolidates the feature details of these datasets, with the ETT dataset being applied in both few-shot and zero-shot learning tasks.
@@
For the short-term forecasting task, we employed the M4 benchmark dataset [Makridakis et al., 2018], which includes 10,000 time series across various domains, from business to economic forecasting, as shown in Table 8. The time series data is categorized into six groups, with sampling rates ranging from annually to hourly.
@@ A.3 BASELINE DETAILS
iTransformer [Liu et al., 2023a]: iTransformer repurposes the Transformer architecture by applying attention and feed-forward networks on inverted dimensions to improve multivariate time series forecasting, achieving state-of-the-art performance on real-world datasets.
@@
FEDformer [Zhou et al., 2022]: FEDformer combines Transformer models with seasonal-trend decomposition to capture both global trends and detailed structures in time series data.
@@
Stationary [Liu et al., 2022]: Stationary introduces Non-stationary Transformers, which include Series Stationarization for improved predictability and De-stationary Attention to restore intrinsic non-stationary information, resulting in significant performance improvements and a substantial reduction in mean squared error compared to mainstream Transformer models.
@@
ETSFormer [Woo et al., 2022]: ETSFormer is a novel Transformer architecture designed specifically for time-series forecasting, addressing the limitations of traditional models by incorporating exponential smoothing principles.
@@
PatchTST [Nie et al., 2022]: PatchTST introduces an efficient Transformer-based architecture for multivariate time series forecasting and self-supervised representation learning, utilizing a segmentation approach that divides time series into subseries-level patches. This design enhances local semantic retention, significantly reduces computation and memory usage of attention maps, and allows the model to consider longer historical data.
@@
TimesNet [Wu et al., 2022]: TimesNet is introduced as a novel approach for time series analysis that focuses on modeling temporal variations by transforming 1D time series into 2D tensors, thereby capturing complex intraperiod and interperiod variations.
@@
Dlinear [Zeng et al., 2023]: Dlinear challenges the efficacy of Transformer-based models for long-term time series forecasting (LTSF), highlighting their limitations in capturing temporal relationships due to the permutation-invariant nature of self-attention mechanisms.
@@
N-HiTs [Challu et al., 2023]: N-HiTs introduces a novel model for long-horizon forecasting that utilizes hierarchical interpolation and multi-rate data sampling techniques to effectively address the challenges of prediction volatility and computational complexity.
@@
N-BEATS [Oreshkin et al., 2019]: N-BEATS addresses the univariate time series point forecasting problem using a novel deep neural architecture featuring backward and forward residual links with a deep stack of fully connected layers. Importantly, the model’s configuration without time-series-specific components suggests that deep learning primitives alone can effectively tackle a variety of forecasting challenges while also providing interpretable outputs with minimal accuracy loss.
@@
GPT4TS [Zhou et al., 2023]: GPT4TS presents a novel approach that leverages pre-trained LLMs for general time series analysis, addressing the challenge of limited training data by utilizing the Frozen Pretrained Transformer (FPT) architecture without altering the self-attention and feedforward layers.
@@
Time-LLM [Jin et al., 2024]: Time-LLM is a reprogramming framework designed to adapt LLMs for general time series forecasting by aligning time series data with natural language through input transformation and context enrichment techniques. Here, we use GPT-2 [Radford et al., 2019] as the base LLM.
@@
S2IP-LLM [Pan et al., 2024]: S2IP-LLM aligns pre-trained LLMs with time series embeddings to enhance forecasting performance. By designing a tokenization module for cross-modality alignment and utilizing semantic anchors from pre-trained word embeddings, S2IP-LLM effectively encodes temporal dynamics and retrieves relevant context for time series.
@@ A.4 Evaluation Metrics
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of long-term, few-shot, and zero-shot forecasting. For short-term forecasting on the M4 benchmark, following the approach in N-BEATS [Oreshkin et al., 2019], we adopt Symmetric Mean Absolute Percentage Error (SMAPE), Mean Absolute Scaled Error (MASE), and Overall Weighted Average (OWA). For time series classification tasks, as referenced in GPT4TS [Zhou et al., 2023], accuracy is used as the evaluation metric.
@@
where $H$ denotes the number of data samples, corresponding to the forecasting horizon in the experiment. $s$ represents the periodicity of the time series. $Y_{h}$ and $\hat{Y}_{h}$ refer to the $h$-th ground truth and its corresponding prediction, respectively.
@@ Appendix B Dual-Scale Context-Alignment GNNs in Classification
The construction of Dual-Scale Context-Alignment GNNs relies on the given prompt. In this section, we introduce the Dual-Scale Context-Alignment GNNs based on the vanilla prompt and few-shot prompting technique in classification tasks. Due to input length constraints, it is challenging to provide examples for many categories. Therefore, for multi-category classification tasks, we employ VCA approach. For binary classification tasks, we use FSCA and select one example from each category to serve as fixed input examples for both training and testing.
@@ B.1 Vanilla Context-Alignment (VCA)
We only replace the prompt content in Sec. 3.2 with “Predict category ($x$ in total) using previous data:”, and the method for constructing the graph remains similar to Sec. 3.2, where $x$ denotes the number of categories.
B.2 Few-Shot prompting Context-Alignment (FSCA)
@@
$\{{\bm{z}}_{1}^{(k)},\dots,{\bm{z}}_{m}^{(k)}\}=\{{\bm{z}}_{1}^{(l+1)},\dots,{\bm{z}}_{m}^{(l+1)}\}=\{{\bm{z}}_{1},\dots,{\bm{z}}_{m}\}$ are the token embeddings of the prompt. We utilise $\{{\bm{e}}_{1}^{(k)},\dots,{\bm{e}}_{n}^{(k)}\}$ as fixed examples during training and testing phases, where $\{{\bm{e}}_{1}^{(k)},\dots,{\bm{e}}_{n}^{(k)}\}$ is from the $k$-th categories and $k=1,\dots,l$. ${\bm{y}}^{(k)}$ is the embedding of the correct label for $\{{\bm{e}}_{1}^{(k)},\dots,{\bm{e}}_{n}^{(k)}\}$. $\{{\bm{e}}_{1}^{(l+1)},\dots,{\bm{e}}_{n}^{(l+1)}\}$ is the TS that needs to be classified. In the fine-grained GNN $G_{F}$, each element in the form 10 is treated as a node. Similar to Sec. 3.2 and Sec. 3.3, two learnable linear layers $f_{e}$, $f_{z}$ map form 10 to form 11. Every element in the form 11 belongs to an $M$-dimensional space.
@@
The directed edges of $G_{F}$ are decompositions of the directed edges of $G_{C}$, which can be represented by the formula LABEL:edgeFC.
@@
Similar to the demo in Sec. 3.2, we introduce an assignment matrix and a learnable weight matrix to achieve learnable interactions between the two scales. Both $\hat{\bm{N}}_{F}$ and $\hat{\bm{N}}_{C}$ are input into pre-trained LLMs. The Dual-Scale Context-Alignment GNNs can be flexibly integrated into various layers of pre-trained LLMs as depicted in Fig. 1, and only the first time apply it needs $f_{e}$ and $f_{z}$ to obtain coarse-grained GNN. The output from the $G_{F}$ branch is used to compute the cross entropy loss against the ground truth.
@@ C.1 Long-term Forecasting full results
Table 10 presents the detailed long-term forecasting results across four prediction horizons. Compared to other models, including LLM-based methods, Transformer-based models, and other high-performing approaches, FSCA demonstrates strong and relatively stable performance across various datasets. It achieves an average 3.1% MSE reduction compared to the second-best method, PatchTST, and outperforms LLM-based methods (S2IP-LLM, Time-LLM, and GPT4TS) by 7.3%, 12.2%, and 16.6%, respectively. We attribute this consistent advantage to the successful Context-Alignment, which effectively guides LLMs’ deep understanding of time series data, with demonstration examples prompt serving as a key factor in enhancement.
@@
width = 1colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,18,23,28,33,381 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18,23,28,33,38 = 1-240.05em, hline43 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21,24-26,29-31,34-36,39-41, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18,23,28,33,38abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22,27,32,37,42belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer